All Parts pleaseExercise 21.49 A charge of -9.00 nC is spread unif
Answer #1

a) The electric field of a disk is:

E_x = 2\pi \frac{q}{\pi R^2} k\left ( 1-\frac{x}{\sqrt{x^2+R^2}} \right )

using given values:

E_x = 2\pi \frac{-9*10^{-9}}{\pi(0.013)^2} 9*10^9\left ( 1-\frac{0.03}{\sqrt{0.03^2+0.013^2}} \right )

E_x = 7.903*10^4 N/C

b) In this case the electric field is given by:

E = \frac{kqx}{(x^2+R^2)^{3/2}} = \frac{9*10^9(9*10^{-9})(0.03)}{(0.03^2+0.013^2)^{3/2}} = 6.95\times 10^{4} N/C

c) In this case the electric field is:

E = \frac{kq}{x^2} = \frac{9*10^{-9}(9*10^9)}{0.03^2} = 90000 N/C